# Inference of multi-channel r-process element enrichment in the Milky Way using binary neutron star merger observations

### v2

@@ Abstract
Observations of GW170817 strongly suggest that binary neutron star (BNS) mergers can produce rapid neutron-capture nucleosynthesis (r-process) elements. However, it remains an open question whether BNS mergers can account for all the r-process element enrichment in the Milky Way’s history. Here we demonstrate that a BNS population model informed by multimessenger neutron star observations predicts a merger rate and per-event r-process element yield consistent with geophysical and astrophysical abundance constraints. If BNS mergers are to explain the r-process enrichment of stars in the Galaxy, we further show using a one-zone Galactic chemical evolution model that they have to merge shortly after the formation of their progenitors, with a delay time distribution of power-law index $\alpha\leq-2.0$ and minimum delay time $t_{\rm min}\leq 40$ Myr at 90% confidence. Such short delay times are in tension with those predicted by standard BNS formation models and those observationally inferred from samples of short gamma-ray bursts (sGRBs). However, we find that a two-channel enrichment scenario, where the second channel follows the star formation history, can account for both Galactic stellar and sGRB observations. Our results suggest that 45–90% of the r-process abundance in the Milky Way today was produced by a star-formation-tracking channel, rather than BNS mergers with significant delay times.
@@ 1 Introduction
A fundamental question of nuclear astrophysics is the origin of the elements in the solar system, the Milky Way, and the Universe as a whole. While it is relatively well-understood how elements up to iron are forged in the cores of stars and ejected through supernovae to enrich the metallicity of future stellar generations, the origins of the rapid neutron-capture (‘r-process’) elements, which must be produced in dense neutron-rich environments, remains open to active discussion (Cowan et al., 2021; Siegel, 2022; Arcones & Thielemann, 2022).
@@
Observational constraints from various astrophysical environments tell us the characteristics of the events that contribute r-process elements to our Universe. Geophysical constraints (Wallner et al., 2015; Hotokezaka et al., 2015) as well as observed levels and scatter of r-process elements in metal-poor stars in different environments, such as the Galactic halo (Macias & Ramirez-Ruiz, 2018), ultra-faint dwarf galaxies (Ji et al., 2016; Beniamini et al., 2016), and globular clusters (Roederer & Sneden, 2011; Roederer et al., 2016), suggest that they must be rare but prolific events in both recent and early Galactic history. For a summary and discussion of rate-yield constraints, see, e.g., Hotokezaka et al. (2018) and Siegel (2019). Observed abundances of r-process tracer elements such as Europium relative to stellar metallicity over the history of iron enrichment by supernovae, from the metal-poor stars of the Galactic halo to the high-metallicity environments of Galactic disk stars, suggest a significant, if not dominant, fraction of Galactic r-process events must have a short delay time with respect to star formation (Matteucci et al., 2014; Wehmeyer et al., 2015; Hotokezaka et al., 2018; Siegel, 2019; Siegel et al., 2019; Côté et al., 2019; Lian et al., 2023; Van Der Swaelmen et al., 2023), although this has been questioned (Banerjee et al., 2020; Tarumi et al., 2021). Furthermore, the brief star formation histories in (ultra-faint) dwarf galaxies of only $\mathcal{O}(100\,\text{Myr})$ and of $\mathcal{O}(10\,\text{Myr})$ in globular clusters also require very short delay times for internal r-process enrichment (Ji et al., 2016; Skúladóttir et al., 2019; Naidu et al., 2022; Zevin et al., 2019; Kirby et al., 2020, 2023).
@@
However, when evaluating candidate channels for r-process contributions in various environments, it has often been assumed that the overall rate of events, the amount of r-process element-containing ejected matter from each event (the ‘yield’), and the delay with respect to star formation are largely unconstrained and can be tuned to the levels required by abundance observations. With the advent of gravitational-wave (GW) astronomy establishing binary neutron star mergers as one site for r-process nucleosynthesis via GW170817, and with subsequently improving constraints on merging neutron star population properties (Abbott et al., 2017a), this is no longer true for neutron star mergers. In this work, we bring together several lines of observational evidence to quantitatively constrain the contribution of binary neutron star (BNS) merger events to r-process nucleosynthesis.
@@
First, we make use of direct constraints on the rates and properties of BNS mergers from LIGO-Virgo-KAGRA (LVK) observations over the first three observing runs (The LIGO Scientific Collaboration et al., 2021; Abbott et al., 2023). Past results (e.g., Hotokezaka et al. (2018); Siegel (2019); Côté et al. (2019)) focus on the rate estimates made immediately after the first observation of a BNS merger GW170817 (Abbott et al., 2017a), which assumed that all BNS mergers have similar properties. Additional events, and (thanks to a well-modeled sensitive volume over the following observing runs) observing time without events, however, furthered our understanding of the merger population as a whole. GW astronomy is now producing rate estimates as a function of component masses, which extend beyond the point-observation of GW170817 (Abbott et al., 2019).
Second, we make use of constraints coming from observations of AT 2017gfo (Abbott et al., 2017b) associated with GW170817, the first direct observation of an r-process enrichment event consistent with a kilonova (Metzger et al., 2010; Metzger, 2020). These observations infer the amount of r-process material produced and thus inform models based on numerical relativity simulations of the merger and postmerger phase to predict the amount of r-process ejected by a given BNS system for a given equation of state (EOS) (Krüger & Foucart, 2020; Nedora et al., 2021). To explain the observed GW170817 kilonova requires taking into account not only ejecta contributions from the dynamical merger phase (Rosswog et al., 1999; Ruffert et al., 1997; Oechslin et al., 2007; Hotokezaka et al., 2013), but also from the postmerger phase in the form of neutrino and magnetically driven winds (Dessart et al., 2008; Perego et al., 2014; Desai et al., 2022; Siegel et al., 2014; Ciolfi et al., 2017; Metzger et al., 2018; Combi & Siegel, 2023; Curtis et al., 2024) from a remnant neutron star and disk winds (Metzger et al., 2008; Fernández & Metzger, 2013; Just et al., 2015; Siegel & Metzger, 2017; Fernández et al., 2019; Combi & Siegel, 2023; Kiuchi et al., 2023) from a postmerger accretion disk (see, e.g., Metzger (2020); Siegel (2019, 2022); Margutti & Chornock (2021); Pian (2021); Radice et al. (2020) for a discussion). As carried out in  Chen et al. (2021); De & Siegel (2021); Siegel (2022), modern simulations predict the contributions through most of these outflows from the component masses of each merger and the neutron-star EOS.
@@
Third, we take a data-driven posterior distribution for the neutron-star EOS at high density, which combines information from an array of astronomical observations of neutron stars. Here, we use the public EOS posterior samples of Legred et al. (2021), which are informed by pulsar mass measurements, GW observations, and X-ray pulse profile modeling. The EOS constraints encoded in this posterior distribution are broadly consistent with those of e.g. Huth et al. (2022); Capano et al. (2020), which include additional astrophysical and nuclear-physics constraints.
@@
Fourth, we take an empirical estimate for the delay time distribution of neutron-star mergers that comes from the observations of the galaxy offsets of short gamma-ray bursts (sGRBs) (Zevin et al., 2022). This provides limits on the distribution of delay times for merger events, assuming that sGRB associations are an unbiased probe of the merging BNS population. Notably, the power-law slope of the sGRB-inferred delay time distribution is steeper than the conventional expectation of $\alpha\approx-1$ for BNSs formed through isolated binary evolution (Piran, 1992; Dominik et al., 2012; Mapelli & Giacobbo, 2018; Neijssel et al., 2019).
@@
Together, these results have quantitative implications for the amount of r-process elements that will be contributed by BNS mergers. We produce four key results:
(1) We estimate the r-process ejecta mass and rate distributions from BNS mergers, as constrained by the observational results mentioned above, and update the comparison of Hotokezaka et al. (2018) and Siegel (2019). We show that the predicted rates and yields overlap with geophysical and astrophysical constraints.
(2) We use a one-zone Galactic chemical evolution model of nucleosynthesis from both supernovae and BNS mergers with the observed rates, masses, EOS, and delay times to derive a distribution of predicted tracks through the relative abundance history of [Fe/H] to [Eu/Fe] (Siegel et al., 2019). As expected, this single-site model does not suffice to generate the amount of Europium enrichment observed at low metallicity. However, as seen in (1), the observed contribution from merger events can quantitatively reproduce the relative abundances observed at solar metallicity. We note that current uncertainties are large, so this does not yet exclude additional significant channels.
(3) We consider what delay-time distribution would be required, instead of the one suggested by sGRBs observations, to reproduce both the Europium abundance at solar metallicity and the changing abundances at lower metallicity. As expected, the significant contribution at low metallicity requires shorter delay times than those inferred from sGRBs.
(4) Starting with the observation-based estimates of the contributions from BNS mergers and retaining the sGRB-inferred delay times, we constrain the overall yield of r-process contribution of a second channel as needed to make up the difference between the contributions of BNS mergers and the observations of high-metallicity disk stars with [Fe/H]$\gtrsim-1$.
@@ 2 Binary neutron star r-process production rate and yield
[…]
@@
    $m_{\rm ej}=\alpha_{\rm dyn}m_{\rm dyn}+f_{\rm loss}m_{\rm disk},$
    where $m_{\rm{dyn}}$ represents the mass of the dynamical ejecta, and $m_{\rm{disk}}$ the mass of the disk formed in the postmerger phase. In order to estimate $m_{\rm{dyn}}$ and $m_{\rm{disk}}$, we use analytical fits to numerical simulations in Krüger & Foucart (2020)—specifically, their Eq. (6) for the dynamical ejecta and Eq. (4) for the disk—which depend on the mass and radius of the component neutron stars. The coefficient $\alpha_{\rm{dyn}}$ is a scaling factor which is randomly sampled between $[0.5,1.5]$ to account for an uncertainty of 50% in the knowledge of $m_{\rm{dyn}}$. The fraction $f_{\rm{loss}}$ of mass ejected from the disk is randomly sampled between $[0.15,1]$ to account for its uncertainty.
Given these three prescriptions, for each rate and EOS sample, we sample uniformly in component masses $(m_{1},m_{2})$ up to the maximum neutron star mass, estimate the ejecta mass for each of these BNS realizations, and perform a rate-density-weighted sum over BNS realizations to obtain the average ejecta mass per event.
In Fig. 1, we show the inferred BNS merger rate as a function of average ejecta mass per event. The r-process production rate and yield for BNS mergers is consistent with constraints from the total stellar mass in the Milky Way (McMillan, 2011), metal-poor stars (Macias & Ramirez-Ruiz, 2018), dwarf galaxies (Beniamini et al., 2016), and ${}^{244}$u measurements (Hotokezaka et al., 2015).
Figure 1: BNS event rate per Milky Way like galaxy as a function of average ejecta mass per event. The dark and light orange shades are the 68% and 90% confidence intervals. We over plot the constraints from total stellar mass in Milky Way ( McMillan (2011), blue band), metal-poor stars ( Macias & Ramirez-Ruiz (2018), purple, lower limit), dwarf galaxies ( Beniamini et al. (2016), grey contours), and ${}^{244}$u measurements ( Hotokezaka et al. (2015), green band). We assume an r-process with the solar abundance pattern for atomic mass number $A\geq 69$.
@@ 3 Chemical abundance evolution
Although the BNS merger rate and average ejecta mass is consistent with existing constraints at $z=0$, a different picture may emerge when enrichment is viewed as a function of cosmic time in various environments and in the context of chemical evolution including other metals. The observed abundances of r-process tracer elements like Eu as a function of stellar metallicity [Fe/H] record the r-process element enrichment at different stages of Milky Way evolution. The r-process enrichment history associated with the BNS merger channel depends on its rate evolution, which can be expressed in terms of a star formation rate convolved with a distribution of delay times measuring the lag between the formation of the progenitor system and the merger. In order to examine whether BNS mergers can produce the measured stellar abundances, we use sGRB observations to inform the BNS delay time distribution and adopt a one-zone model to estimate chemical abundances over time.
@@
    $R_{\rm sGRB}(t)$ $=$
[…]
    where $\Psi_{\rm SFH}(t)$ represents the Galactic star formation rate per unit time; we assume the Galactic star formation history follows the cosmic star formation history (Madau & Fragos, 2017). The normalization factor $C_{\rm sGRB}$ is calibrated by the observed local sGRB rate. The delay time distribution (DTD)
@@
  • One-zone model: We describe the Galactic production of iron-peak and r-process elements over cosmic time using the one-zone chemical evolution model of Siegel et al. (2019). One-zone models assume instantaneous homogeneous mixing of metals in the interstellar medium (ISM), and thus cannot adequately capture effects of hierarchical structure growth and inhomogeneities due to incomplete mixing associated with rare enrichment events at low metallicity. However, they can accurately capture the average enrichment levels of the ISM with r-process elements at high metallicity, once the ISM has been polluted by a significant number of individual events. We use the same Galactic in- and outflow rate and the same rates and yields for iron production by core-collapse and Type Ia supernovae as in Siegel et al. (2019), while we adjust the rate and yield of r-process enrichment as traced by Eu according to the local merger rate, DTD, and ejecta mass estimates discussed above. The total r-process ejecta mass per event is translated into a Eu mass per event by assuming a solar abundance distribution of r-process elements (Arnould et al., 2007) starting at atomic mass number $A=69$. Representing the cumulative total over many individual enrichment events and a rough average to the r-process abundance patterns of many individual observations of r-process enhanced metal-poor stars that have been polluted by a single or only by a few r-process events(e.g. Frebel & Ji 2023), the solar r-process abundance pattern provides a reasonable estimate for average r-process enrichment at high metallicity.
@@
We first consider an r-process enrichment scenario in which BNS mergers with an sGRB-informed DTD are the only enrichment site. Fig. 2 shows the abundance ratio of [Eu/Fe] as a function of [Fe/H]. We overplot the observed metallicities of Galactic disk stars from Battistini & Bensby (2016) (blue diamonds), as well as observations from the Stellar Abundances for Galactic Archaeology (SAGA) database (Suda et al. (2011); green crosses), which span both metal-rich disk stars and metal-poor halo stars. Despite the large statistical uncertainty originating from the uncertainties in BNS merger rate (locally and at higher redshift), mass distribution, neutron star EOS, and analytic fits to ejecta in numerical simulations as described in the previous section, the abundance ratio [Eu/Fe] estimated from the BNS-only model shows a clear deviation from the stellar observations at low metallicity. Thus, while this single-site model can match the observed Eu abundance at $z=0$, it cannot reproduce the Galactic history of r-process enrichment.

### v3

@@ Abstract
Observations of GW170817 strongly suggest that binary neutron star (BNS) mergers produce rapid neutron-capture nucleosynthesis (r-process) elements. However, it remains an open question whether these mergers can account for all the r-process element enrichment in the Milky Way’s history. Here, we constrain the contributions of the BNS channel using astrophysical neutron star observations. The rate and mass distributions are constrained by LIGO/Virgo/Kagra through the latest catalog GWTC-3, the neutron star equation of state by gravitational-wave, radio, and X-ray observations, and the delay time distribution by short gamma-ray burst (GRB) host galaxy associations. We present a Bayesian framework to consistently combine these observations with abundance information to quantify the contribution and uncertainties of single and multiple astrophysical enrichment sources, and obtain a distribution of per-event BNS r-process element yields consistent with geophysical and astrophysical abundance constraints. We then adopt a galactic chemical evolution model assuming instantaneous and fixed amount of Fe enrichment from core-collapse supernovae, and show that BNS-only enrichment scenarios remain inconsistent with the observed r-process abundance trend of disk stars in the Galaxy even with the uncertainties in BNS merger observations. Using stellar abundance observations instead of the short GRB constraints, we can infer a shorter BNS delay time distribution with power-law index $\alpha\leq-2.0$ and minimum delay time $t_{\rm min}\leq 40$ Myr at 90% confidence, consistent with detailed galactic chemical evolution models. Such delay times are in tension with those predicted by standard BNS formation models. Alternatively, we confirm that a two-channel scenario, in which the second channel tracks the star formation history without significant delay, can account for both Galactic stellar and short GRB observations. We estimate that 45–90% of the r-process abundance in the Milky Way today would have been produced by this star-formation-tracking channel, rather than BNS mergers with significant delay times.
@@ 1 Introduction
A fundamental question of nuclear astrophysics is the origin of the elements in the solar system, the Milky Way, and the Universe as a whole. While it is understood how elements up to iron are forged in the cores of stars and ejected through supernovae to enrich the metallicity of future stellar generations, the origins of the rapid neutron-capture (‘r-process’) elements, which must be produced in dense neutron-rich environments, remains open to active discussion (Cowan et al., 2021; Siegel, 2022; Arcones & Thielemann, 2022).
@@
Observational constraints from various astrophysical environments tell us the characteristics of the events that contribute r-process elements to our Universe. Geophysical constraints (Wallner et al., 2015; Hotokezaka et al., 2015) as well as observed levels and scatter of r-process elements in metal-poor stars in different environments, such as the Galactic halo (Macias & Ramirez-Ruiz, 2018), ultra-faint dwarf galaxies (Ji et al., 2016; Beniamini et al., 2016), and globular clusters (Roederer & Sneden, 2011; Roederer et al., 2016), suggest that they must be rare but prolific events in both recent and early Galactic history—see, e.g., Hotokezaka et al. (2018) and Siegel (2019). Observed abundances of r-process tracer elements such as Europium relative to stellar metallicity over the history of iron enrichment by supernovae, from the metal-poor stars of the Galactic halo to the high-metallicity environments of Galactic disk stars, suggest a significant fraction of Galactic r-process events must have a short delay time with respect to star formation (Matteucci et al., 2014; Wehmeyer et al., 2015; Hotokezaka et al., 2018; Siegel, 2019; Siegel et al., 2019; Côté et al., 2019; Kobayashi et al., 2020; Lian et al., 2023; Van Der Swaelmen et al., 2023), although this has been questioned (Banerjee et al., 2020; Tarumi et al., 2021). Furthermore, the brief star formation histories in (ultra-faint) dwarf galaxies of only $\mathcal{O}(100\,\text{Myr})$ and of $\mathcal{O}(10\,\text{Myr})$ in globular clusters also require very short delay times for internal r-process enrichment (Ji et al., 2016; Skúladóttir et al., 2019; Naidu et al., 2022; Zevin et al., 2019; Kirby et al., 2020, 2023).
@@
However, when evaluating candidate channels for r-process contributions in various environments, it has often been assumed that the overall rate of events, the amount of r-process element-containing ejected matter from each event (the ‘yield’), and the delay with respect to star formation are largely unconstrained and can be tuned to the levels required by abundance observations. In addition, previous work exploring multi-channel r-process enrichment (Matteucci et al., 2014; Wehmeyer et al., 2015; Côté et al., 2019; Kobayashi et al., 2020, 2023) focused on point estimates for fixed scenarios, such as a fixed binary neutron star (BNS) merger rate or a fixed fractional secondary channel contribution. The advent of gravitational-wave (GW) astronomy establishing BNS mergers as one site for r-process nucleosynthesis via GW170817, the subsequently improving constraints on merging neutron star population properties (Abbott et al., 2017), on the equation of state of NSs as well as on the delay time distribution of sGRBs provide a wide array of empirical information that strongly affects the modeling of chemical evolution of r-process elements. In this work, we bring together several new lines of observational evidence and develop a Bayesian inference framework to appropriately take into account the uncertainties in different observational products to quantitatively constrain the contribution of BNS merger events to r-process nucleosynthesis.
@@
First, we make use of direct constraints on the rates and properties of BNS mergers from LIGO-Virgo-KAGRA (LVK) observations over the first three observing runs (The LIGO Scientific Collaboration et al., 2021; Abbott et al., 2023). Past results (e.g., Hotokezaka et al. (2018); Siegel (2019); Côté et al. (2019)) focus on the rate estimates made immediately after the first observation of a BNS merger GW170817 (Abbott et al., 2017), which assumed that all BNS mergers have similar properties. Additional events, and (thanks to a well-modeled sensitive volume over the following observing runs) observing time without events, however, furthered our understanding of the merger population as a whole. GW astronomy is now producing rate estimates as a function of component masses, which extend beyond the point-observation of GW170817 (Abbott et al., 2019).
@@
Second, we take a data-driven posterior distribution for the neutron-star EOS at high density, which combines information from an array of astronomical observations of neutron stars. Here, we use the public EOS posterior samples of Legred et al. (2021), which are informed by pulsar mass measurements, GW observations, and X-ray pulse profile modeling. The EOS constraints encoded in this posterior distribution are broadly consistent with those of e.g. Huth et al. (2022); Capano et al. (2020), which include additional astrophysical and nuclear-physics constraints.
@@
Third, we take an empirical estimate for the delay time distribution of neutron-star mergers that comes from the observations of the galaxy offsets of short gamma-ray bursts (sGRBs) (Zevin et al., 2022). This provides limits on the distribution of delay times for merger events, assuming that sGRB associations are an unbiased probe of the merging BNS population. Notably, the power-law slope of the sGRB-inferred delay time distribution is steeper than the conventional expectation of $\alpha\approx-1$ for BNSs formed through isolated binary evolution (Piran, 1992; Dominik et al., 2012; Mapelli & Giacobbo, 2018; Neijssel et al., 2019).
@@
Together, these new observations have quantitative implications for the amount of r-process elements that will be contributed by BNS mergers. We introduce our methods and present the four key results in Sections 2 to  5. We then summarize and discuss the results in Section 6.
@@ 2 Binary neutron star r-process production rate and yield
  • Local merger rate: We use the BNS population distribution inferred from the LVK’s GWTC-3 catalog for the local ($z<0.1$) merger rate (The LIGO Scientific Collaboration et al., 2021). Specifically, we adopt the Power Law + Dip + Break population model from Abbott et al. (2023); other population model choices in that study predict consistent merger rates. Abbott et al. (2023) provides 1501 posterior rate samples binned in mass. For every rate sample, we sum the rate density over mass bins up to the maximum mass for a given neutron star EOS sample and convert the resulting BNS merger rate to a Galactic event rate, $R_{\rm MW}$, assuming a number density of Milky-Way like galaxies of 0.01 Mpc-3 (Gehrels et al., 2016).
@@
    $m_{\rm ej}=\alpha_{\rm dyn}m_{\rm dyn}+f_{\rm loss}m_{\rm disk},$ (1)
    where $m_{\rm{dyn}}$ represents the mass of the dynamical ejecta, and $m_{\rm{disk}}$ the mass of the disk formed in the postmerger phase. In order to estimate $m_{\rm{dyn}}$ and $m_{\rm{disk}}$, we use analytical fits to numerical simulations in Krüger & Foucart (2020)—specifically, their Eq. (6) for the dynamical ejecta and Eq. (4) for the disk—which depend on the mass and radius of the component neutron stars. The coefficient $\alpha_{\rm{dyn}}$ is a scaling factor which is randomly sampled between $[0.5,1.5]$ to account for an uncertainty of 50% in the knowledge of $m_{\rm{dyn}}$. The fraction $f_{\rm{loss}}$ of mass ejected from the disk is randomly sampled between $[0.15,1]$ to account for its uncertainty. Both of these uncertainties are taken from the variations in existing numerical simulations of neutron-star mergers (Chen et al., 2021).
Given these three prescriptions, for each rate and EOS sample, we sample uniformly in component masses $(m_{1},m_{2})$ up to the maximum neutron star mass, estimate the ejecta mass for each of these BNS realizations, and perform a rate-density-weighted sum over BNS realizations to obtain the average ejecta mass per event. In Fig. 1, we show the inferred BNS merger rate as a function of average ejecta mass per event.
We compare our results to different geographical and astrophysical constraints summarized in  Hotokezaka et al. (2018), and find them consistent with the following constraints from observations:
  • The total galactic r-process mass: Assuming the stellar mass in the Milky Way of $6.4\times 10^{10}\,M_{\odot}$ (McMillan, 2011) follows a solar abundance pattern and integrating the mass of elements with atomic number $A\geq 69$ to obtain the total mass of r-process elements. The average ejecta mass per event is then a function of r-process event rate assuming the age of the Milky Way is $\sim 10$ Gyr.
  • Lower limit on the r-process yield per event: The limit comes from abundance observations of metal-poor stars that have presumably been enriched by a single event only (Macias & Ramirez-Ruiz, 2018).
  • Dwarf galaxies: The difference between the observations of strongly disparate r-process enrichment in dwarf and ultra-faint dwarf galaxies indicates the rarity of r-process events, which can be converted into rate (Beniamini et al., 2016).
  • Geophysical 244Pu measurements in the deep sea crust and meteorites: The deep sea crust measurements represent r-process deposition over the past tens of Myrs, and those from meteorites indicate the abundance $\approx\!4.5$ Gyr ago. Their total abundance and the comparison between the two jointly constrain the rate and yield of r-process events (Hotokezaka et al., 2015)
We note that various assumptions had to be made to produce the geographical and astrophysical constraints summarized above (see  Hotokezaka et al. (2018) for more details). The consistency between our results only suggests that BNS mergers can explain the rate and yield derived from other observations, but it does not exclude the possibility of other r-process enrichment channels.
Figure 1: BNS event rate per Milky Way like galaxy as a function of average ejecta mass per event. The dark and light orange shades are the 68% and 90% confidence intervals. We over plot the constraints from total stellar mass in Milky Way ( McMillan (2011), blue band), metal-poor stars ( Macias & Ramirez-Ruiz (2018), purple, lower limit), dwarf galaxies ( Beniamini et al. (2016), grey contours), and 244Pu measurements ( Hotokezaka et al. (2015), green band). We assume an r-process with the solar abundance pattern for atomic mass number $A\geq 69$.
@@ 3 Chemical abundance evolution
Although the BNS merger rate and average ejecta mass is consistent with existing constraints at high metallicity, where $z\simeq 0$, a different picture may emerge when enrichment is viewed as a function of cosmic time in various environments and in the context of chemical evolution including other metals. The observed abundances of r-process tracer elements like Eu as a function of stellar metallicity [Fe/H] record the r-process element enrichment at different stages of Milky Way evolution (Matteucci, 2021). The r-process enrichment history associated with the BNS merger channel depends on its rate evolution, which can be expressed in terms of a star formation rate convolved with a distribution of delay times measuring the lag between the formation of the progenitor system and the merger. In order to examine whether BNS mergers can produce the measured stellar abundances, we use sGRB observations to inform the BNS delay time distribution and adopt a simplified one-zone model from Siegel et al. (2019) to estimate chemical abundances over time.
@@
    $R_{\rm sGRB}(t)$ $=$ (2)
    $C_{\rm sGRB}$ $\int_{0}^{t}D(t-t^{\prime}|t_{\rm min},t_{\rm max},\alpha)\Psi_{\rm SFH}(t^{\prime})dt^{\prime},$
    where $\Psi_{\rm SFH}(t)$ represents the Galactic star formation rate per unit time; we assume the Galactic star formation history follows the cosmic star formation history (Madau & Fragos, 2017). The normalization factor $C_{\rm sGRB}$ represents the observed local sGRB rate. The delay time distribution (DTD)
@@
  • One-zone model: We describe the Galactic production of iron-peak and r-process elements over cosmic time using the one-zone chemical evolution model of Siegel et al. (2019). One-zone models assume instantaneous homogeneous mixing of metals in the interstellar medium (ISM), and thus cannot adequately capture effects of hierarchical structure growth and inhomogeneities due to incomplete mixing associated with rare enrichment events at low metallicity. However, they can accurately capture the average enrichment levels of the ISM with r-process elements at high metallicity, once the ISM has been polluted by a significant number of individual events. We use the same Galactic in- and outflow rate and the same observationally inferred rates and yields for iron production by core-collapse and Type Ia supernovae as in Siegel et al. (2019). Iron-producing core-collapse supernovae are assumed to follow star formation without delay with a local rate of $7.05\times 10^{-5}$ Mpc-3 yr-1 (Li et al., 2011) and a yield of 0.074 $M_{\odot}$ per event (Maoz & Graur, 2017). Type Ia supernovae produce 0.7 $M_{\odot}$ iron per event and, consistent with that assumption, they are assumed to follow the star formation history with a DTD of the form Eq. (3) with $\alpha=-1.0$ and $t_{\rm min}=40$ Myr (Maoz & Graur, 2017), calibrated to a production efficiency of $1.3\times 10^{-3}$ per $M_{\odot}$ of stellar mass formed (Maoz & Graur, 2017). We adjust the rate and yield of r-process enrichment as traced by Eu according to the local merger rate, DTD, and ejecta mass estimates discussed above. The total r-process ejecta mass per event is translated into a Eu mass per event by assuming a solar abundance distribution of r-process elements (Arnould et al., 2007) starting at atomic mass number $A=69$. Representing the cumulative total over many individual enrichment events and a rough average to the r-process abundance patterns of many individual observations of r-process enhanced metal-poor stars that have been polluted by a single or only by a few r-process events (e.g. Frebel & Ji 2023), the solar r-process abundance pattern provides a reasonable estimate for average r-process enrichment at high metallicity. We note that this assumption was made due to the limit of detailed r-process abundance pattern from BNS mergers, and deviations from the solar pattern have been shown (Kobayashi et al., 2023).
@@
We first consider an r-process enrichment scenario in which BNS mergers with an sGRB-informed DTD are the only enrichment site. Fig. 2 shows the abundance ratio of [Eu/Fe] as a function of [Fe/H]. We overplot the observed metallicities of Galactic disk stars from Battistini & Bensby (2016) (blue diamonds), as well as observations from the Stellar Abundances for Galactic Archaeology (SAGA) database (Suda et al. (2011); green crosses), which span both metal-rich disk stars and metal-poor halo stars. Despite the large statistical uncertainty originating from the uncertainties in BNS merger rate (locally and at higher redshift), mass distribution, neutron star EOS, and analytic fits to ejecta in numerical simulations as described in the previous section, the abundance ratio [Eu/Fe] estimated from the BNS-only model shows a clear deviation from the stellar observations at low metallicity, consistent with what was found in previous work. Thus, while this single-site model can match the average observed Eu abundance at solar metallicity, it cannot reproduce the Galactic history of r-process enrichment.